# Anytime Replanning of Robot Coverage Paths for Partially Unknown Environments

### v1

@@ Abstract
In this paper, we propose a method to replan coverage paths for a robot operating in an environment with initially unknown static obstacles. Existing coverage approaches reduce coverage time by covering along the minimum number of coverage lines (straight-line paths). However, recomputing such paths online can be computationally expensive resulting in robot stoppages that increase coverage time. A naive alternative is greedy detour replanning, i.e., replanning with minimum deviation from the initial path, which is efficient to compute but may result in unnecessary detours. In this work, we propose an anytime coverage replanning approach named OARP-Replan that performs near-optimal replans to an interrupted coverage path within a given time budget. We do this by solving linear relaxations of mixed-integer linear programs (MILPs) to identify sections of the interrupted path that can be optimally replanned within the time budget. We validate our approach in simulation using maps of real-world environments and compare our approach against a greedy detour replanner and other state-of-the-art approaches.
@@ I Introduction
One approach to coverage replanning is to stop executing the current coverage path and plan a new path for the uncovered environment with the new obstacle(s). However, coverage planning is an NP-Hard problem [5] and this approach does not scale well for large environments with many unknown obstacles. Moreover, an obstacle may only affect a small section of the path, i.e., the unbroken parts of the path may still be drivable and near-optimal. An alternate approach is to only replan the affected section(s) of the path (local replanning) while keeping the rest of the path intact. A simple example of local replanning would be where the robot simply drives around the obstacle whenever a collision is imminent. However, a more practical approach would be to use the time while driving towards the blockage to find more optimal solutions. In this paper, we propose an anytime coverage replanner that utilizes the time spent driving to conduct near-optimal replans on the provided coverage paths (see Fig. 1).
@@
    We present a Mixed-Integer Program (MILP) that replans the ranks of a coverage path such that a tour of the ranks (i.e. a coverage path) can be replanned within a time budget.
@@
      We leverage LP relaxations of this MILP to replan ranks in polynomial time (efficient to compute).
@@
      We present a second MILP that reduces the number of decision variables, and therefore replans the coverage ranks faster than the first MILP.
@@
    We provide an anytime coverage replanning method named OARP-Replan that leverages the above MILPs to conduct near-optimal replans to the coverage path within a given time budget.
@@
    Finally, we conduct simulations using maps of real-world environments that show that our proposed approach achieves better coverage performance than other state-of-the-art approaches, including an offline approach (i.e. all obstacles are known) [7].
@@ I-A Related Works
Grid-based approaches use a grid approximation of the given environment and compute a tour of the grid cells to compute a coverage path. This tour can be computed either by solving the Traveling Salesman Problem (TSP) [8], using a spanning tree of the grid cells (Spanning Tree Coverage or STC) [15], or using a wavefront algorithm to determine visitation order [16]. Turn-minimizing approaches of this class look to minimize the number of axis-parallel (horizontal or vertical) coverage lines or ranks in the coverage path [17, 18]. While these methods were demonstrated to work well in offline settings, they lack online extensions that can replan a path within a given time budget. Our proposed method, OARP-Replan, extends the OARP approach from [6], which is an offline grid-based turn-minimizing approach, to be applicable in online replanning scenarios.
@@
Online Coverage Planning: Online coverage planning approaches have gained recent interest, and as such, there is a wide range of different approaches in the literature. Initially, online exact decomposition approaches were investigated [19, 20], but these methods involve identifying changes to exact cells online, which does not scale well for large cluttered environments. Moreover, these methods do not aim to minimize turns in the coverage path, while our proposed method does.
@@
Grid-based methods work better for such environments as one can easily update a grid with new occupancy information. Our proposed approach falls under this class. A popular approach of this class is Online STC [21], which computes a spanning tree of the grid cells online to update its coverage plan. However, this method does not minimize turns. Other grid-based online approaches use various local search and optimization methods to determine the coverage plan in an unknown environment. Song et al [22] proposed an approach using artificial potential fields (APF) to determine a coverage path online. A similar technique to APFs is using bio-inspired neural networks to model and search the environment. This was originally proposed by Luo and Yang in [23] and has gained popularity in many recent works [24, 25, 26]. However, these local search-based approaches are susceptible to deadlocking, where the robot does not know where to go next. Deadlocks are which is computationally expensive to get out of and affect overall coverage time. Moreover, these approaches generally minimize turns by penalizing them in a local search cost function, which does not guarantee minimum turns for the path overall. Our proposed approach avoids deadlocks (using an initial path) and guarantees minimum turns in the overall path when provided with sufficient planning time.
Other recent online approaches include decomposition-free planning [27]. However, it still remains to be shown whether such methods can effectively plan paths that optimize coverage time. Online coverage replanning has also been investigated in 3D settings for applications in underwater inspection tasks [28], but they involve 3D maneuvers and are not necessarily feasible for robots operating in 2D planes.
@@ I-B Organization
The rest of this paper is organized as follows. In Section II, we provide some background on the coverage planning problem and the OARP approach from our previous work [6]. In Section III, we define the problem of replanning a coverage path within a given time budget. In Section IV, we present the high-level framework of our proposed approach, OARP-Replan, that constrains the runtime to replan the coverage path, which solves the problem in two stages: (i) rank replanning, and (ii) touring. In Section V, we describe the rank replanning stage where we use MILPs to replan the ranks of the coverage path such that the overall runtime is constrained. In Section VI, we describe the touring stage where we formulate a Generalized Traveling Salesman Problem (GTSP) to compute a tour of the replanned ranks to obtain the replanned path. In Section VII, we present experimental results on a range of real-world environments and compare our proposed approach against state-of-the-art coverage planning methods [6, 7] in both online and offline settings.
@@ II Preliminaries
In this section, we review the coverage path planning problem and our previously proposed coverage planning approach, namely the optimal axis-parallel rank partitioning (OARP) approach [6].
@@ II-A The Coverage Planning Problem
Consider a 2D non-convex environment with known obstacles, where $\mathcal{W}\subseteq\mathbb{R}^{2}$ is a closed and bounded set that represents all free points within the environment. We now consider a robot with a state composed of its position $x\in\mathcal{W}$ and a heading $\theta\in[0,2\pi)$ radians. The robot is carrying a coverage tool with a footprint given by $\mathcal{A}\subseteq\mathbb{R}^{2}$. In this work, similar to [29, 17, 6], we assume that the coverage tool is a square of width $l>0$. Let $\mathcal{A}(x,\theta)\subseteq\mathcal{W}$ represent the placement of the tool with respect to the robot’s state. In reality, a square tool may only be able to cover a subset of the environment $\widetilde{\mathcal{W}}\subseteq\mathcal{W}$. However, for simplicity of notation, we assume that $\widetilde{\mathcal{W}}=\mathcal{W}$. We now formally define the coverage planning problem (CPP) of interest as follows.
@@
Given a 2D environment $\mathcal{W}$ and a robot carrying a coverage tool of footprint $\mathcal{A}$, plan a path $P$ from the set $\mathcal{P}$ of all possible paths which solves
@@ II-B Optimal Axis-Parallel Rank Partitioning (OARP)
We now briefly describe the coverage planning approach from our previous work [6], namely the Optimal Axis-Parallel Rank Partitioning (OARP) approach. The focus of the OARP approach is to plan optimal coverage paths in which the coverage tool moves only in axis-parallel directions (horizontal and vertical) during coverage. The OARP approach uses a three-step framework to plan coverage paths: (i) decompose the environment $\mathcal{W}$ into an Integral Orthogonal Polygon (IOP), which is a set of square grid cells of size $l$, (ii) place the minimum number of axis-parallel (horizontal or vertical) ranks (straight-line paths) to cover the IOP, and (iii) compute a tour of the ranks to construct the coverage path. To place the minimum number of ranks in the IOP, OARP solves a linear program (LP) to compute the coverage orientation of each cell in the IOP. Once the ranks are obtained, a tour of the ranks is computed by solving a Generalized Travelling Salesman Problem (GTSP) that minimizes the cost of transitioning (driving) between ranks. The resulting coverage path is represented as a series of ranks connected by cost-minimizing transition paths, which are collision-free dynamically feasible paths to traverse between ranks. Note that, unlike ranks, transition paths do not have to be axis-parallel.
@@ III-A Coverage Replanning Problem
Let us consider the case where our robot is trying to follow its initial coverage path $P$ but it observes a set $\mathcal{O}$ of previously unknown obstacles. In other words, the actual environment is $\bar{\mathcal{W}}\subseteq\mathcal{W}\setminus\mathcal{O}$. We assume the robot is equipped with a sensor that can observe its environment within a sensor footprint relative to the robot’s position, e.g., an optical range sensor. Suppose, at position $x_{i}$, the robot detects a new obstacle interrupting the path. Let $\tau$ be the time in which the robot would reach the obstacle along its current path from $x_{i}$. While the robot might be able to continue along $P$ for time $\tau$, the subsequent path must be replanned. One solution is to solve the full CPP for the new environment $\bar{\mathcal{W}}$. While this method would give us the optimal coverage path, it has a major drawback. If solving CPP takes longer than $\tau$, the robot may have to stop and wait for a new path, which is undesirable. To address this, we instead solve the coverage replanning problem that takes into account a time budget $\tau$.
Let $P_{\text{r}}$ be the remaining path after the first interruption by the obstacle. Let $\mathcal{W}_{\text{r}}\subseteq\mathcal{W}$ be the area covered by $P_{\text{r}}$, where $\mathcal{O}\subseteq\mathcal{W}_{\text{r}}$. The replanning problem is to plan a new coverage path $P_{\text{r}}^{\prime}$ before $\tau$ such that it covers the same areas as $P_{\text{r}}$ without colliding into $\mathcal{O}$. We are now ready to pose the main problem we look to solve in this paper.
@@
[…]
@@ III-B Rank-based Coverage Replanning
[…]
@@ IV-B Algorithm
Figure 3: An example run of OARP-Replan with one obstacle. Each encounter (blue dots) with the new obstacles along the current path is evaluated. For encounter $i\leq k$, the planner attempts to replan the remaining path (red path) within the time $\tau_{i}$ to reach the encounter (black path). To compute $\tau_{i}$, previous encounters are replanned using GD replan (green dotted lines).
@@
We now describe the high-level algorithm of our proposed approach, namely OARP-Replan, for which the pseudocode is given in Algorithm 1. OARP-Replan is run when new obstacles blocking the current path $P$ are detected. Let $k$ be the number of projected encounters with all newly detected obstacles along $P$ (each obstacle may cause multiple encounters). For each encounter, the robot has the option of replanning the remaining path (the preferred solution) or simply going around the obstacle while covering the interrupted rank (if there is no time to replan). We refer to the latter method as the greedy detour (GD) replan, which can be performed using a local planner with little to no computational time and thus no robot stoppage. We propose only using GD replan as a fallback method in the case where the main planner does not have enough time to replan the path.
@@
Fig. 3 illustrates an example run of the algorithm. We iterate through each obstacle encounter and evaluate whether the main planner can replan the path after the $i^{\text{th}}$ encounter, where $i\leq k$. In other words, we look for an obstacle encounter that gives the planner enough time to replan the path after the encounter, i.e,. the path for all previous obstacle encounters ($\{1,2,\dots i-1\}$) is replanned using GD replan. Let $\tau_{i}$ be the time for the robot to reach the $i^{\text{th}}$ encounter along the current path, which will be the time budget for the replan.
@@
To evaluate if the path after encounter $i$ can be replanned, the planner runs a sub-routine named RANK-REPLAN to get a set of replanned ranks for which a tour can be computed in time $\tau_{i}$ (Line 9). If RANK-REPLAN returns a set of replanned ranks, we compute a tour of the ranks using another sub-routine called TOUR-REPLAN (Line 11) that solves the GTSP to obtain our replanned path. This sub-routine runs as a background process while the robot continues to follow the current coverage path. However, if RANK-REPLAN returns an empty set, then this indicates that there is no solution that will finish within the expected time budget. In that case, GD replan is used to replan encounter $i$ (Line 14), and the procedure is repeated for the next encounter. In the end, we obtain a replanned path where some (or all) encounters are replanned using GD replan, while the remaining path undergoes a near-optimal replan before the robot reaches the corresponding encounter.
@@
1:Current path $P$, Environment $\mathcal{W}$, New obstacles $\mathcal{O}$
@@
3:$k\leftarrow$ No. of encounters with $\mathcal{O}$ along $P$
4:for $i=1,2,\dots,k$ do
5:     $\mathcal{W}_{i}\leftarrow$ Part of $\mathcal{W}$ to cover after $i^{\text{th}}$ encounter
6:     $R\leftarrow$ Ranks from coverage path $P$
7:     $\tau_{i}\leftarrow$ Time to reach $i^{\text{th}}$ encounter
8:     $\bar{m}_{i}\leftarrow\hat{T}^{-1}(\tau_{i})$triangleright Maximum instance size given $\tau_{i}$
9:     $R^{\prime}\leftarrow$ RANK-REPLAN($\mathcal{W}_{i}$, $R$, $\bar{m}_{i}$)
10:     if $R^{\prime}\neq\emptyset$ then
11:         $P^{\prime}\leftarrow$ TOUR-REPLAN($R^{\prime}$, $R$)
12:         return $P^{\prime}$
13:     else
14:         Apply GD replanning to $i^{th}$ encounter

### v2

@@ Abstract
In this paper, we propose a method to replan coverage paths for a robot operating in an environment with initially unknown static obstacles. Existing coverage approaches reduce coverage time by covering along the minimum number of coverage lines (straight-line paths). However, recomputing such paths online can be computationally expensive resulting in robot stoppages that increase coverage time. A naive alternative is greedy detour replanning, i.e., replanning with minimum deviation from the initial path, which is efficient to compute but may result in unnecessary detours. In this work, we propose an anytime coverage replanning approach named OARP-Replan that performs near-optimal replans to an interrupted coverage path within a given time budget. We do this by solving linear relaxations of integer linear programs (ILPs) to identify sections of the interrupted path that can be optimally replanned within the time budget. We validate OARP-Replan in simulation and perform comparisons against a greedy detour replanner and other state-of-the-art coverage planners. We also demonstrate OARP-Replan in experiments using an industrial-level autonomous robot.
@@ I Introduction
One approach to coverage replanning is to stop executing the current coverage path and plan a new path for the uncovered environment with the new obstacle(s). However, coverage planning for arbitrary environments is generally NP-Hard [5] and this approach does not scale well for large environments with many unknown obstacles. Moreover, an obstacle may only affect a small section of the path, i.e., the unbroken parts of the path may still be drivable and near-optimal. An alternate approach is to only replan the affected section(s) of the path (local replanning) while keeping the rest of the path intact. A simple example of local replanning would be where the robot simply drives around the obstacle whenever a collision is imminent. However, a more practical approach would be to use the time while driving towards the blockage to find more optimal solutions. In this paper, we propose an anytime coverage replanner that utilizes the time spent driving to conduct near-optimal replans on the provided coverage paths (see Fig. 1).
@@
    We present an Integer Linear Program (ILP) that replans the ranks of a coverage path such that a tour of the ranks (i.e. a coverage path) can be replanned within a time budget.
@@
      We leverage LP relaxations of this ILP to replan ranks in polynomial time (efficient to compute).
@@
      We present a second ILP that reduces the number of decision variables, and therefore replans the coverage ranks faster than the first ILP.
@@
    We provide an anytime coverage replanning method named OARP-Replan that leverages the above ILPs to conduct near-optimal replans to the coverage path within a given time budget.
@@
    Finally, we conduct simulations and robot experiments to validate our proposed approach. Our simulation results include comparisons against other state-of-the-art approaches, including offline approaches (i.e. all obstacles are known) [7, 8].
@@ I-A Related Works
Grid-based approaches use a grid approximation of the given environment and compute a path covering all the grid cells in the approximation. This path can be computed either by solving the Traveling Salesman Problem (TSP) [9], using a spanning tree of the grid cells (Spanning Tree Coverage or STC) [16], or using a wavefront algorithm to determine visitation order [17]. Turn-minimizing approaches of this class look to minimize the number of axis-parallel (horizontal or vertical) coverage lines or ranks in the coverage path [18, 19]. While these methods were demonstrated to work well in offline settings, they lack online extensions that can replan a path within a given time budget. Our proposed method, OARP-Replan, extends the OARP approach from [6], which is an offline grid-based turn-minimizing approach, to be applicable in online replanning scenarios.
@@
Online Coverage Planning: Online coverage planning approaches have gained recent interest, and as such, there is a wide range of different approaches in the literature. Initially, online exact decomposition approaches were investigated [20, 21], but these methods involve identifying changes to exact cells online, which does not scale well for large cluttered environments. Moreover, these methods do not aim to minimize turns in the coverage path.
@@
Grid-based methods work better for such environments as one can easily update a grid with new occupancy information. Our proposed approach falls under this class. A popular approach of this class is Online STC [22], which computes a spanning tree of the grid cells online to update its coverage plan. However, given a square coverage tool of width $l$, STC is generally applied to coarse grids with cells of width $2l$ and does not generalize well to finer grids with cell width $l$. Other grid-based online approaches use various local search and optimization methods to determine the coverage plan in an unknown environment. Song et al [23] proposed an approach using artificial potential fields (APF) to determine a coverage path online. A similar technique to APFs is using bio-inspired neural networks to model and search the environment. This was originally proposed by Luo and Yang in [24] and has gained popularity in many recent works [25, 26, 27]. However, these local search-based approaches are susceptible to deadlocking, where the robot does not know where to go next. Deadlocks are computationally expensive to get out of and affect overall coverage time. Moreover, these approaches generally minimize turns by penalizing them in a local search cost function, which does not guarantee minimum turns for the path overall. Our proposed approach avoids deadlocks (using an initial path) and guarantees minimum turns in the overall path when provided with sufficient planning time.
Online coverage planning shares similarities with exploration planning, which involves covering unknown environments using sensors [28]. Exploration planning algorithms use information theory [29, 30] and frontiers [31] to determine the next-best scan of the environment, rather than using decompositions. Recently, a decomposition-free online coverage planner was proposed that uses frontiers and search-based planning to identify the next-best region to cover [32]. Online coverage replanning has also been investigated in 3D settings for applications in underwater inspection tasks [33], but they involve 3D maneuvers and are not necessarily feasible for robots operating in 2D planes.
@@ I-B Organization
The rest of this paper is organized as follows. In Section II, we provide some background on the coverage planning problem and the OARP approach from our previous work [6]. In Section III, we define the problem of replanning a coverage path within a given time budget. In Section IV, we present the high-level framework of our proposed approach, OARP-Replan, that constrains the runtime to replan the coverage path, which solves the problem in two stages: (i) rank replanning, and (ii) touring. In Section V, we describe the rank replanning stage where we use ILPs to replan the ranks of the coverage path such that the overall runtime is constrained. In Section VI, we describe the touring stage where we formulate a Generalized Traveling Salesman Problem (GTSP) to compute a tour of the replanned ranks to obtain the replanned path. In Section VII, we provide simulation results on a range of real-world environments and compare our proposed approach against state-of-the-art coverage planning methods in online and offline settings. In Section VIII, we present experimental results of testing our replanning approach using an industrial-level autonomous cleaning robot.
@@ II Preliminaries
We first review the coverage planning problem and our previously proposed coverage planning approach, namely the optimal axis-parallel rank partitioning (OARP) approach [6].
@@ II-A The Coverage Planning Problem
Consider a 2D non-convex environment with known obstacles, where $\mathcal{W}\subseteq\mathbb{R}^{2}$ is a closed and bounded set that represents all free points within the environment. We now consider a robot with a state composed of its position $\boldsymbol{x}\in\mathcal{W}$, where $\boldsymbol{x}$ is a vector in $\mathbb{R}^{2}$, and a heading $\theta\in[0,2\pi)$ radians. The robot is carrying a coverage tool with a footprint given by $\mathcal{A}\subseteq\mathbb{R}^{2}$. In this work, similar to [34, 18, 6], we assume that the coverage tool is a square of width $l>0$. Let $\mathcal{A}(\boldsymbol{x},\theta)\subseteq\mathcal{W}$ represent the placement of the tool with respect to the robot’s state. In reality, a square tool may not be able to cover the entire environment, so we look to cover an approximation of the environment $\widetilde{\mathcal{W}}\subseteq\mathcal{W}$ that the tool can reach. Consider a set $\mathcal{P}$ of all possible paths in the environment, where each path $P\in\mathcal{P}$ is given by a sequence of poses, i.e. position and heading pairs $(\boldsymbol{x},\theta)$. We now formally define the coverage planning problem (CPP) of interest as follows.
@@
Given an approximation $\widetilde{\mathcal{W}}$ of a 2D environment $\mathcal{W}$ and a robot carrying a coverage tool of footprint $\mathcal{A}$, plan a path $P$ from the set $\mathcal{P}$ of all possible paths which solves
@@ II-B Optimal Axis-Parallel Rank Partitioning (OARP)
We now briefly describe the coverage planning approach from our previous work [6], namely the Optimal Axis-Parallel Rank Partitioning (OARP) approach. The focus of the OARP approach is to plan optimal coverage paths in which the coverage tool moves only in axis-parallel directions (horizontal and vertical) during coverage. The OARP approach uses a three-step framework to plan coverage paths: (i) decompose the environment $\mathcal{W}$ into an Integral Orthogonal Polygon (IOP), which is a set of square grid cells of size $l$, (ii) place the minimum number of axis-parallel (horizontal or vertical) ranks (straight-line paths) to cover the IOP, and (iii) compute a tour of the ranks to construct the coverage path. Here, the IOP forms the environment subset $\widetilde{\mathcal{W}}$ that we look to cover. To place the minimum number of ranks in the IOP, OARP solves a linear program (LP) to compute the coverage orientation of each cell in the IOP. Once the ranks are obtained, a tour of the ranks is computed by solving a Generalized Travelling Salesman Problem (GTSP) that minimizes the cost of transitioning (driving) between ranks. The resulting coverage path is represented as a series of ranks connected by cost-minimizing transition paths, which are collision-free dynamically feasible paths to traverse between ranks. Note that, unlike ranks, transition paths do not have to be axis-parallel.
@@ III-A Coverage Replanning Problem
Let us consider the case where our robot is trying to follow its initial coverage path $P$ but it observes a set $\mathcal{O}$ of previously unknown obstacles. In other words, the actual environment is $\overline{\mathcal{W}}\subseteq\mathcal{W}\setminus\mathcal{O}$. We assume the robot is equipped with a sensor that can observe its environment within a sensor footprint relative to the robot’s position, e.g., an optical range sensor. Suppose, at position $\boldsymbol{x}_{i}$, the robot detects a new obstacle interrupting the path. Let $\tau$ be the time in which the robot would reach the obstacle along its current path from $\boldsymbol{x}_{i}$. While the robot might be able to continue along $P$ for time $\tau$, the subsequent path must be replanned. One solution is to solve the full CPP for the new environment $\overline{\mathcal{W}}$. While this method would give us the optimal coverage path, it has a major drawback. If solving CPP takes longer than $\tau$, the robot may have to stop and wait for a new path, which is undesirable. To address this, we instead solve the coverage replanning problem that takes into account a time budget $\tau$.
Let $P_{\text{r}}$ be the remaining path after the first interruption by the obstacle. Let $\mathcal{W}_{\text{r}}\subseteq\widetilde{\mathcal{W}}$ be the area covered by $P_{\text{r}}$, where $\mathcal{O}\cap\mathcal{W}_{\text{r}}\neq\emptyset$. The replanning problem is to plan a new coverage path $P_{\text{r}}^{\prime}$ before $\tau$ such that it covers the same areas as $P_{\text{r}}$ without colliding into $\mathcal{O}$. We are now ready to pose the main problem we look to solve in this paper.
@@
Given a path $P$ covering $\widetilde{\mathcal{W}}$, a set of newly observed obstacles $\mathcal{O}$, a time $\tau$ until the first such obstacle is encountered along $P$, the remaining path $P_{\text{r}}$ after the first encounter, and the area $\mathcal{W}_{\text{r}}\subseteq\widetilde{\mathcal{W}}$ covered by $P_{\text{r}}$, find a new coverage path $P_{\text{r}}^{\prime}$ within time $\tau$ which solves
$\min_{P_{\text{r}}^{\prime}\in\mathcal{P}}\quad J(P_{\text{r}}^{\prime})$ (3)
@@ III-B Rank-based Coverage Replanning
Given a set of newly observed obstacles $\mathcal{O}$ that interrupts the remainder $P_{\text{r}}$ of the coverage path $P$, and a time budget $\tau$ before the robot’s first encounter with an obstacle in $\mathcal{O}$, compute the following within time $\tau$: (i) a minimum set of ranks that covers the IOP representation of the uncovered free region $\mathcal{W}_{\text{r}}\setminus\mathcal{O}$ and (ii) a tour of the ranks that minimize the total cost of transition between ranks.
@@ IV-B Algorithm
Figure 3: A run of OARP-Replan with one obstacle (dotted grey box). The visible part of the obstacle (pink dots) is used to identify blocked grid cells (grey) and obstacle encounters (blue dots) along the current path. For each encounter $i\leq k$, the planner attempts to replan the remaining path (red path) within the time $\tau_{i}$ to reach the encounter (black path). To compute $\tau_{i}$, previous encounters are replanned using GD replan (green dotted lines).
@@
We now describe the high-level algorithm of our proposed approach, namely OARP-Replan, for which the pseudocode is given in Algorithm 1. We consider a robot that is equipped with a sensor similar to that of a laser scanner (LIDAR) that is capable of detecting obstacles (or parts of obstacles) in the environment. OARP-Replan is run when potential collisions or encounters are detected along the current path $P$ using the robot’s sensor model. More specifically, we determine the grid cells in the IOP decomposition that cannot be covered anymore due to newly detected obstacles (if the obstacle was partially observed, this process would be repeated for each scan that reveals more of the obstacle). Let $k$ be the number of all projected encounters along $P$ (each obstacle may cause multiple encounters). For each encounter, the robot has the option of replanning the remaining path (the preferred solution) or simply going around the obstacle while covering the interrupted rank (if there is no time to replan). We refer to the latter method as the greedy detour (GD) replan, which can be performed using a local planner with little to no computational time and thus no robot stoppage. We propose only using GD replan as a fallback method in the case where the main planner does not have enough time to replan the path.
@@
Fig. 3 illustrates an example run of the algorithm. Note that we replan using the obstacle encounters that have been observed so far by the robot. The encounters that were not observed (even from an obstacle that was partially observed) is replanned once the robot has observed them later in the coverage run. We iterate through each obstacle encounter $i\leq k$ and evaluate whether the main planner can replan the path after the $i^{\text{th}}$ encounter. The path for all previous obstacle encounters ($\{1,2,\dots i-1\}$) is replanned using GD replan. In other words, we look for an obstacle encounter that gives the planner enough time to replan the path after the encounter. Let $\tau_{i}$ be the time for the robot to reach the $i^{\text{th}}$ encounter along the current path, which will be the time budget for the replan.
@@
To evaluate if the path after encounter $i$ can be replanned, the planner runs a sub-routine named RANK-REPLAN to get a set of replanned ranks for which a tour can be computed in time $\tau_{i}$ (Line 8). If RANK-REPLAN returns a set of ranks, we compute a tour of the ranks using another sub-routine called TOUR-REPLAN (Line 10) that solves the GTSP to obtain our replanned path. This sub-routine runs as a background process while the robot continues to follow the current coverage path. However, if RANK-REPLAN returns an empty set, then this indicates that there is no solution that will finish within the expected time budget. In that case, GD replan is used to replan encounter $i$ (Line 13), and the procedure is repeated for the next encounter. In the end, we obtain a replanned path where some (or all) encounters are replanned using GD replan, while the remaining path undergoes a near-optimal replan before the robot reaches the corresponding encounter.
@@
1:Current path $P$, Environment IOP $\widetilde{\mathcal{W}}$
@@
3:for each encounter $i\leq k$ on $P$ do
4:     $\mathcal{W}_{i}\leftarrow$ Part of $\widetilde{\mathcal{W}}$ to cover after $i^{\text{th}}$ encounter
5:     $R\leftarrow$ Ranks from coverage path $P$
6:     $\tau_{i}\leftarrow$ Time to reach $i^{\text{th}}$ encounter
7:     $\bar{m}_{i}\leftarrow\hat{T}^{-1}(\tau_{i})$triangleright Maximum instance size given $\tau_{i}$
8:     $R^{\prime}\leftarrow$ RANK-REPLAN($\mathcal{W}_{i}$, $R$, $\bar{m}_{i}$)
9:     if $R^{\prime}\neq\emptyset$ then
10:         $P^{\prime}\leftarrow$ TOUR-REPLAN($R^{\prime}$, $R$)
11:         return $P^{\prime}$
12:     else
13:         Apply GD replanning to $i^{th}$ encounter